Status of Assembled Hit Finder Boards
Last updated on $Date: 2000/07/17 00:43:12 $ (UTC)
-
49 good, 0 bad total rev B boards, out of 42 needed (including L00), leaving 7 hot and 0 cold spares
-
2 rev B boards good after SMT rework
-
27 rev B boards good first try
-
16 rev B boards good after we fixed minor problems
-
4 rev B boards good after Twin Hunter replaced BGA chips
-
0 rev B board still not working
-
2 rev A boards working, with some kludges, different firmware, etc.
-
Board 01 - March 1999 prototype
-
Assembled by ADCO; arrived 1999-03-24
-
Has broken DAD pin; requires "pinswap" version of DAD program as a work-around
-
Otherwise, no known problems
-
Originally, hf_vme_addr bit 11 tracked bit 14; traced problem down to HF (not boot) side of HCT244 buffer (U106), presumably pins 9,16; was never able to demonstrate problem with ohm meter, but WJA fixed 1999-03-29 by removing and resoldering U106
-
Originally, grt_data bit 94 (highest data bit of stream 9) was stuck at 0; did not see any DC short with ohm meter; problem between P3-D47 and U92-230; broke pin U92-230 while trying to unsolder it; replaced with U92-240 as a kludge (WJA 1999-06-04)
-
Currently at Fermilab as Vertical Slice Test data source
-
Board 02 - March 1999 prototype
-
Assembled by ADCO; arrived 1999-03-24
-
Originally, grt_data bit 98 (clock input for G-link #2) was stuck; presumably bad connection on U92-211; Mircea fixed 1999-04-14 by cleaning U92 with alcohol
-
Currently at Chicago as Track Fitter data source
-
Board 03 - Jan 2000 pre-production
-
Assembled by SMT; arrived 2000-01-26; reworked twice (arrived 2000-04-20 and 2000-05-03); installed at Fermilab 2000-05-05
-
Summary: after two passes of SMT rework, board functions perfectly, and is installed at Fermilab
-
No power shorts; power-up; power LED OK
-
Both EPC2 and Boot chip seen by JTAG programmer
-
Boot chip programmed directly via JTAG
-
Load mode LED on; error LED flashing to indicate unconfigured Flex chain, OK
-
Able to read brain register via VME; DTACK flashes; change mode to load/run/test; see LEDs
-
Configure EPC2; EPC2 correctly programs Boot chip
-
Load FRAM OK; Boot chip programs Flex chain OK; program Merger chip via JTAG OK
-
Temporary confusion due to pilot error, described below
-
Can W/R random contents to CRAM, ISPY, PedRAM, misc registers, OK
-
Can send test patterns from GSTM through to ISPY; all data bits are OK for all streams from P3 to DAD to HitMan to ISPY
-
FIFO W/R detects problems with stream 4 (bit 9), stream 8 (FIFO is always empty), stream 9 (bits 0-9) (out of 18 total bits/stream, 10 streams) (problems on streams 8 and 9 are now fixed; problem on stream 4 remains)
-
OSPY W/R detects that bit 16 is stuck off, and bit 7 is floating
-
problem with bit 7 seems to be between MOP and OSPY, as bit 7 works for I/O to other MOP registers
-
problem with bit 16 seems to be between MOP and HF_VME_Data bus, as other VME reads from MOP have same problem, but VME reads from e.g. ISPY do not show problems with this bit
-
Except for the bad bits, cluster data are correctly processed and merged, until stream 8's permanently empty FIFO stalls the pipeline
-
2000-02-04: On stream 8 FIFO (U65), clock/control signals were seen on both sides of FIFO with scope; pin 19 (wclk) was bent; several pins appeared not to be soldered; Mircea resoldered; problem fixed
-
On stream 9 FIFO (U59), many pins in a row were not soldered; Mircea resoldered; problem fixed
-
Stream 4 FIFO input bit 9 is floating, because HM4 (U37) pin L22 (fifodata9) appears not to be soldered to the BGA; we verified that trace is connected to FIFO; scope does not see activity on trace when BGA output is driven; JTAG scan does not see change in state when voltage is applied to trace
-
MOP VME data bit 16 appears to be a missing solder connection on MOP (U30) pin P22, as JTAG does not see stimuli applied to this bit via TC9-19 (but does see stimuli on TC9-18, e.g.)
-
MOP OSPY data bit 7 appears to be a missing solder connection on MOP (U30) pin J22, as JTAG does not see stimuli applied to this trace, but trace is clearly connected to OSPY (U7) pin 59.
-
Went through list of PCB changes (prototype vs production) on 2000-02-01, and found everything to be OK
-
All JTAG chains have been used to detect/identify/configure their respective chips
-
Flex configuration works, so pullup resistors on CONF_DONE and nSTATUS are evidently doing their job
-
GSTM data move successfully into ISPY buffers, so HitMan dadclock input clearly works; hf_init was shown to be working by looking at FIFO reset signal on a scope; vmeaddr0-2 and vmeclock functionality demonstrated by W/R test to pedestal RAM (load entire memory with random constants and read it back)
-
FRAM contents OK after 5V power on with 3V fuse removed, which was reason for pulldown on RP* line; check with ohm-meter confirms presence of pull-up/down resistors (on RP* and VPEN); voltages have desired values when 3V fuse is removed
-
Able to program FRAM, so its VPEN voltage is OK now
-
Power LED now requires both 3.3V and 5V power to be present; checked by removing each fuse in turn
-
LEDs (e.g. HOLD light) now light when they should, not when they shouldn't; in particular HOLD LEDs behave properly
-
Front panel fits comfortably, so LED mount positions are fine
-
Signals sent from MOP & Boot chips for DS, HOLD, CdfError, SvtError LEDs were inverted, because that was what had been required to make them work with kludges on prototypes--now fixed in firmware
-
Initially loaded old (prototype) version of HitMan program, which caused many temporary problems that I chose not to describe above--correct version now loaded
-
back from SMT rework 2000-04-20
-
undid change that made hitman E2,F2 dummy input pins (didn't affect this board anyway)
-
data path from all HM chips to FIFOs to Merger to MOP is fine; problem with stream 4 FIFO is fixed
-
MOP connections to OSPY and VME bus are fine; OSPY random fill/read now works
-
looping 50-event burst/check test using ISPY as data source works (650 loops OK and counting); but doesn't work from GSTM (new problem) for streams 1 and 5
-
I see that HM1,5 (U35,U36) think that data valid is always asserted; also true coming out of DAD (TC3-15); not true when I unconfigure the chips; need to track down with boundary-scan
-
News flash! Mircea discovers that about 40 pins on U92 (a DAD chip) are not soldered to the board!
-
back from second SMT rework 2000-05-03
-
Board 04 - Jan 2000 pre-production
-
Assembled by SMT; arrived 2000-01-26; reworked board arrived 2000-04-20; installed at Fermilab 2000-05-02
-
Summary: after SMT rework, board functions perfectly, and is installed at Fermilab
-
Initially, +3.3V was shorted to GND; Mircea and Harold traced the problem to a solder-blob short beneath the upper-right corner of U33 (HM3) and cleaned out the solder blob
-
SMT agreed to use a better technology (different flux, e.g.) when assembling remaining boards
-
Merger chip programmed OK
-
EPC2 and Boot chip seen via JTAG; EPC2 programmed OK; EPC2 loads Boot OK; VME I/O to Boot chip OK; program FRAM OK
-
Auto-boot of 13-chip Flex chain fails, even though FRAM contents are correct; need to use JTAG to diagnose, narrow down to a particular chip (now fixed, see below)
-
2000-02-02: With scope, I see that CONF_DONE is coming up about 14 clock cycles before end of second dad.ttf (end of Flex chain), indicating successful configuration, but that about 5 usec thereafter, nCONFIG is pulled back down for 150 nsec, after which the chips return to their unconfigured state
-
Hypothesis is that one of the 13 chips in the Flex chain (probably a BGA) has an output shorted to nCONFIG
-
By loading "empty" 10K50 firmware (no logic; all pins are inputs) into MOP and HitMen, I was able to get board to configure, consistent with hypothesis above
-
Narrowed problem down to HM8; can load normal firmware into other twelve chips in Flex chain
-
Now time for a few tests on the chips that are working:
-
misctest OK; ospytest OK; ispytest OK; pramtest OK; cramtest OK; fifotest OK except stream 7 bit 14; board ID switches OK; dadtest OK
-
2000-02-03: Both ohm-meter and JTAG boundary scan indicate that HM8 (U38) has a short between pin D2 (nCONFIG) and pin E2 (an unused pin, marked RESERVED in the compiler report file)
-
First use of boundary scan was misleading, because I got the sign wrong and was looking at chip two instead of chip four in a five-chip chain
-
Changing pin E2 to a dummy INPUT pin is a successful work-around for the short beneath HM8
-
But stream 7 FIFO (U61) bit 14 still stuck at GND (now fixed, see below)
-
Four possible locations: output of HM7, input of FIFO, output of FIFO, input of Merger; last is ruled out by other working streams
-
See activity on scope on both input (pin 2) and output (pin 48) of FIFO; notice that "fifotest" on stream 7 succeeds when scope probe is on pin 48
-
Problem is fixed by heating pin 48 with iron
-
Board now processes data successfully, 48 events at a time, in burst/check mode, for ~ 5000 iterations (1.5 hours)
-
Connected each output connector in turn to TF; checked data in TF ISPY; no problems
-
Ran random test (burst/check) overnight for ~ 15 hours, 40K loops of 48 events, checking both HF OSPY and TF ISPY; no problems
-
2000-02-06: Am pumping data continuously from GSTM, polling OSPY with C code on crate CPU
-
GSTM loops over 50 events; 12686 words in each FIFO, of which 63% have DAV asserted
-
12686 words / 55 MHz = 230 usec per loop, or 4350 loops/sec, or 217K events/sec
-
HF output is 5706 ospy words/loop, 24M words/sec
-
polling program checks about 64K words about 7.5 times/sec, or 490K words/sec, or 2% of total HF throughput
-
this is much more throughput than HF would see in real beam data
-
100KB of SIXD (seems like a plausible upper bound; do we have a hard number?) shared between 36 hit finders at a 50 kHz event rate would put DAV up ~ 25% of the time, compared with 63% in the test
-
Guram's "B10" svtsim sample gave an estimate of 36 clusters per HF per event, which would be about 2M words/sec at 50 kHz event rate, compared with 24M words/sec in the test
-
in ~ 21 hours, I checked 36.8 x 10^9 ospy words, no errors
-
back from SMT rework 2000-04-20
-
powers up, boots, processes data in burst/check mode from GSTM
-
undid change that made hitman E2,F2 dummy input pins--still OK!
-
processed 1700 loops of 50-event burst/check test--OK!
-
I think the board is fine
-
Board 05 - Apr 2000 pre-production ("first item")
-
Assembled by Twin Hunter; arrived 2000-04-28; installed at Fermilab 2000-05-02
-
Summary: Board is perfect, and is installed at Fermilab
-
Mircea checked for shorts, powered up, configured EPC2
-
VME I/O then worked right away; Bill loaded FRAM while Mircea configured Merger chip
-
All single-board VME tests passed
-
Burst/check test from GSTM has run for ~ 21 hours without failing: 73K loops, 3.5M total events
-
I see that CDF error and SVT error LEDs have on/off inverted: problem was that older version of boot program had been loaded (fix was put in 2000-01-29); now correct program is loaded
-
Interesting note: it seems that boot program compiled with Altera v9.5 does not correctly configure 13-chip Flex chain, while boot program compiled with altera v9.21 (what we were using before) is fine
-
Both FP outputs OK to TF; all LEDs OK
-
Continuous test from GSTM, polling OSPY, checked 6.54 x 10^10 OSPY words over weekend, no errors.
-
Board 06 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Standard set of tests is this:
-
configure boot, merger; configure FRAM; check ID prom
-
run ispy data flow test 100 loops, then 20 loops to TF per connector; check LEDs
-
run gstm burst 100 loops; run gstm continuous 5 minutes (150M ospy words) on crate cpu
-
Weird things occasionally seen:
-
ispy test mode can be screwed up at some low rate by presence of vtm? (so I use non-vtm slot for this test)
-
boards sometimes need reseating in slots
-
Board 07 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
bfifoio bit 15 was stuck on; fixed solder bridge at U63-50/49 (Bill)
-
Board 08 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 09 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 10 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 11 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
delayed_modsel_ OK; delayed_DS_ OK; AM1_ OK; configuration via backup port works (can talk to single register on each chip), but after I turn "backup" mode off, chips become unconfigured
-
removed solder bridge LED3-1/3 (Mircea)
-
Clearing FRAM lock bit before FRAM erase fixes write problem
-
GSTM test passes (1500 loops) if I disable stream 1 in test
-
Back from Twin Hunter rework (replaced HM1); board did not respond to VME i/o, until delay line U81 was replaced. Now passes all GSTM tests.
-
Board 12 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 13 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Fixed solder bridge on U92-200/201 (Mircea)
-
Board 14 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 15 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 16 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 17 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
hm_data(117), connecting U92-87 to U35-A15, is shorted to ground (1 ohm with meter); suspect beneath BGA, where pin A14 is ground; removing debris beneath U35 (Mircea) fixed problem
-
Board 18 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
hf_vme_data, bits 13 and 15, are driven low, apparently by U39 (HM2); no short found with meter
-
after Twin Hunter rework (replaced HM2), board passes all tests
-
Board 19 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 20 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 21 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
wrong LED (G/G) on lower hold/DS replaced with R/G (Mark)
-
Board 22 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
removed a number of hairline solder bridges (Bill) from top row of U91 (even caused ispy burst tests to fail occasionally!)
-
Board 23 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 24 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Fixed solder bridge on U92-65/66 (Mircea)
-
Board 25 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
replaced U89 (Mark) because output did not change in response to input on VME address lines 18-21
-
fixed solder bridge U17-31/32 (Bill); replaced U105 (Bill), which unfortunately resulted in a ruined pad, requiring a jumper
-
U22 is upside down; fixed by Mark
-
Board 26 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 27 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 28 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 29 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 30 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Replaced U107 (Mark); fixed solder bridge U13-31/32 (Mircea); fixed cold solder joints U68-18/20 (Mircea)
-
Board 31 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 32 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 33 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 34 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
ispy 0 bit 9 was stuck at 0; fixed unsoldered pin U18-8 (Bill)
-
Board 35 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
wrong LED (G/G) on upper hold/DS replaced with R/G (Mark)
-
Board 36 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 37 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Fixed solder bridge U61-94/95 (Mircea)
-
Board 38 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 39 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 40 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Fixed solder bridge on U13-45/44 (Bill)
-
Board 41 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Broken trace hm_vme_data(10), just left of U37, jumpered (Mircea/Mark); solder bridge U92-41/42, removed (Mircea)
-
Board 42 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Fixed two solder bridges on U91, at approximately pin 95 (Bill)
-
Board 43 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
suspect HM3 is bad, but not at all certain
-
disabling stream 3 brings ~ 100% failure rate on GSTM test down to ~ 0.03% (2/6000), now on stream 8 (disabling 8 also gives 0 errors after 11600 loops)
-
low speed gstm test works now for ~ 200 loops, but high speed test fails
-
second rework (replaced HM8) done 2000-07-13; GSTM tests now pass!
-
Board 44 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
all non-GSTM tests pass, GSTM tests fail, always on stream 5, but ispy data look fine; HM5 dadclock looks fine
-
Replaced U81 (Mark), since delayed_modsel_ was dead (fixes bus error on block transfer); clearing FRAM lock bit before FRAM erase fixes write problem
-
With stream 5 disabled, GSTM test passes (1650 loops)
-
Back from Twin Hunter rework; now passes GSTM tests! (June 23)
-
Board 45 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Fixed U61-2,4,5 unsoldered pins (Bill)
-
Board 46 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 47 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Replaced capacitors C16,17 (Mircea), but it was a red herring; fixed short U91-179/180, which was making boot fail (and changing effective capacitance); removed solder bridges (Mircea) U91-225/.../230
-
Board 48 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 49 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 50 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests
-
Board 51 - May 2000 production by Twin Hunter (arrived 2000-05-17)
-
Summary: passes all tests